# Inland-LOAM: Voxel-Based Structural Semantic Mapping for Inland Waterways

### v1

@@ 3.3.1 Preprocessing and Plane Fitting
Given the LiDAR’s non-uniform vertical distribution of its 32 laser beams and a fixed horizontal resolution of $0.2^{\circ}$, each scanning cycle produces a 2D range image, $I_{k}$, of dimension $32\times 1800$ (rows $\times$ columns). Each row corresponds to an individual laser beam, $v_{j}$ represents the scan line index (row), while each column represents sequential horizontal angle increments, $u_{j}$ represents the horizontal angle index (column). The resulting range image is shown in Fig. 4.
Figure 4: The single-frame point cloud is preprocessed with the vessel docked alongside the shore. In the upper figure, different laser beams are colored according to their vertical angles, and the red bounding box indicates the physical dimensions of the vessel. Several scan lines are observed to intersect the vessel, and these points are subsequently identified and removed. The lower figure shows the result after preprocessing. The water surface points are represented by blue, and the non-water surface points by yellow.
@@
Our system is designed for short-range localization and mapping, the curvature of the Earth can be neglected, and the water surface can be approximated as a plane [47]. Moreover, considering the limitations discussed above posed by the incidence angles of grazing on the surface of the water, the candidate points for the estimation of the plane are restricted to a predefined subset of lower LiDAR beams ($N_{\text{lines}}$), determined by the installation height of the LiDAR installation and the draft of the vessel.
@@
$\mathbf{n}_{s}^{\top}\mathbf{p}+d_{s}=0$ (2)
where $\mathbf{n}_{s}=[a_{s},b_{s},c_{s}]^{\top}\in S^{2}$ is the normal vector of the plane and $d_{s}\in\mathbb{R}$ is its scalar offset from the sensor origin. Subsequently, any point $\mathbf{p}_{j}\in\mathbf{P}_{k}^{valid}$ is classified as a water surface point, setting the corresponding $M_{k}^{ground}(v_{j},u_{j})=1$, if its perpendicular distance to this estimated plane, calculated as $|\mathbf{n}_{s}^{\top}\mathbf{p}_{j}+d_{s}|$, is within a predefined tolerance $\tau_{s}$.
@@
All points belonging to a connected component found via BFS are assigned a unique positive integer identifier in Channel 3 ($M_{k}^{\text{label}}$). To ensure the robustness of subsequent feature extraction, a final filtering step is applied: segments comprising fewer than 30 points are considered insignificant or noise. Points belonging to these small segments are relabeled in Channel 3 as outliers. Fig 4 illustrates the result after preprocessing.
3.3.2 Feature Extraction
@@
Traditional approaches often extract features based on local range differences. For instance, a common metric (Equation 3, where $|S|$ is the size of the local neighborhood, e.g., 10 points, and $r_{i}$ is the range of point $i$) is:
$c=\frac{1}{|S|\cdot\|r_{i}\|}\left\|\sum_{j\in S,j\neq i}(r_{j}-r_{i})\right\|$ (3)
While this method is computationally efficient and applicable to all directions, it demonstrates limitations in inland waterway environments. In areas with dense vegetation, grass, or gravel, the local surface exhibits numerous small, “edge-like” bumps. When applying Equation 3, the positive and negative range residuals $(r_{j}-r_{i})$ from neighboring points $p_{j}$ on either side of the central point $p_{i}$ tend to cancel each other out. Consequently, such curvature estimations often overlook subtle surface textures and primarily respond only to more extreme geometric features like significant peaks or occlusions.
To overcome this limitation and effectively capture local surface roughness, thereby distinguishing between smooth and textured surfaces, we employ a variance-based approach to quantify local range variations as detailed in Algorithm  1. Fig. 5 illustrates the comparative performance of feature extraction between LOAM-based methods and our optimized approach.
@@
3.3.3 LiDAR Odometry, Mapping and Joint Optimization
Our LiDAR odometry and mapping parts build upon LEGO-LOAM [15], with specific adaptations in Joint Optimization. The system employs a two-stage registration process for pose estimation, integrating water surface constraints. The overall factor graph structure is illustrated in Fig. 6.
Two-Stage LiDAR Odometry: a two-step Levenberg-Marquardt optimization to estimate the transformation between consecutive scans. First, the planar features $\mathcal{F}_{p}^{t}$ extracted from the water surface and other planar environmental elements are matched to their correspondences in $\mathcal{F}_{p}^{t-1}$ to determine $t_{z},\theta_{roll},\theta_{pitch}]$. Then, the remaining transformation parameters $t_{x},t_{y},\theta_{yaw}]$ are estimated using edge features $\mathcal{F}_{e}^{t}$ matched against $\mathcal{F}_{e}^{t-1}$, while applying the previously computed vertical and angular parameters as constraints, resulting in the full 6 DoF transformation. This decoupled registration method significantly reduces computation time and is well-suited to our edge computing scenario.
@@ 3.3.3 LiDAR Odometry, Mapping and Joint Optimization
Joint Optimization with Water Surface Constraints: To mitigate drift, we extend the pose graph optimization framework to incorporate global water surface constraints. We parameterize the water surface as a globally consistent plane $P_{w}$ with parameters $\{\mathbf{n}_{w},d_{w}\}$ as defined in Equation 2.
@@
In this formulation, $\|r\|^{2}_{\Sigma}$ denotes the squared Mahalanobis distance $r^{T}\Sigma^{-1}r$, where $\Sigma$ is the covariance matrix of the corresponding residual $r$. The term $r_{0}$ represents a combined prior residual on initial knowledge about the system, such as the initial pose $T_{0}$ or an initial estimate for the plane, with $\Sigma_{0}$ as its covariance. The odometry residual, $r_{\mathrm{odom},k}$, for the relative motion $\Delta\widetilde{T}_{k,k+1}$ measured between consecutive poses $T_{k}$ and $T_{k+1}$, is given by
@@
where $\mathrm{Log}(\cdot)$ maps an $SE(3)$ element to its tangent space representation. Similarly, the loop closure residual, $r_{\mathrm{loop},ij}$, for a measured relative transform $\Delta\widetilde{T}_{i,j}$ between non-consecutive poses $T_{i}$ and $T_{j}$ (where $(i,j)\in\mathcal{L}$ denotes the set of loop closures), is defined as
@@
$(R_{m}n_{s,m})^{T}x_{w}+\bigl{(}d_{s,m}-(R_{m}n_{s,m})^{T}t_{m}\bigr{)}=0,$ (9)
@@
$\begin{split}r_{\mathrm{plane},m}(T_{m},P_{w})&=P_{w}\ominus\widehat{P}_{w,m}\\ &=\begin{pmatrix}e_{N}\!\bigl{(}n_{w},\hat{n}_{w,m}\bigr{)}\\ d_{w}-\hat{d}_{w,m}\end{pmatrix}\\ &=\begin{pmatrix}e_{N}\!\bigl{(}n_{w},R_{m}n_{s,m}\bigr{)}\\ d_{w}-\bigl{(}d_{s,m}-(R_{m}n_{s,m})^{\mskip 2.0mu\top}t_{m}\bigr{)}\end{pmatrix}\in\mathbb{R}^{3}.\end{split}$ (11)
@@
$\hat{\xi}=-\frac{\arccos(n_{w}^{T}\hat{n}_{w,m})}{\sqrt{1-(n_{w}^{T}\hat{n}_{w,m})^{2}}}\,(\hat{n}_{w,m}-(n_{w}^{T}\hat{n}_{w,m})\,n_{w}),$ (12)
@@
By leveraging the water surface as a global geometric constraint, our approach reduces vertical drift without requiring external references such as GNSS. Fig. 7 demonstrates how this approach effectively maintains a consistent water surface representation throughout the trajectory.
@@
3.4 Structural Semantic Map Conversion
3.4.1 Semantic voxel transformation
In the preceding section, a map comprising structured point cloud features was generated. Each point in this cloud has been classified with semantic labels, including water surface, flat points, or rough points. To efficiently store and query this semantic information within the global map, the ”dense” semantically annotated point cloud is transformed into a ”sparse” voxel grid representation. For each voxel, a multi-class semantic confidence accumulation is maintained, based on the principles of probabilistic occupancy grids [43].
The workspace is discretized into a 3D voxel grid with a uniform resolution $r$, set to $0.15\text{ m}$ in this work. This resolution is chosen considering a balance between memory consumption, computational efficiency, and the typical scale of objects of interest in inland waterways. The entire transformation process is depicted in Fig. 8. For each voxel $v_{i}$, we maintain a hash map to store its semantic state $S(v_{i})$. This state involves accumulating evidence for the semantic classes: water surface, flat terrain, and rough terrain. The ’unknown’ category is assigned if a voxel has insufficient evidence for any specific class.
@@ 3.4.1 Semantic voxel transformation
Points pre-classified as water surface by the LiDAR odometry module are handled directly, if such a point falls into voxel $v_{i}$, the voxel is immediately assigned the ’water’ semantic label with high confidence.
@@
$S(v_{i})=\begin{cases}\text{'flat'}&\text{if }P_{i}^{\text{flat}}-0.5>\tau_{\text{flat}}\\ \text{'rough'}&\text{if }P_{i}^{\text{flat}}-0.5<-\tau_{\text{rough}}\\ \text{'unknown'}&\text{otherwise}\end{cases}$
3.4.2 2D structural semantic map construction
The 3D semantic voxel map is subsequently projected onto a unified 2D structural semantic map. While it refers on the conversion method [40], this variant targets USV rather than UGV/UAV scenarios and adopts different map architectures for refined slope estimation and richer semantic inference. To effectively map the complex environment encountered by USVs in inland waterways, this framework broadly classifies non water surface areas into four structural semantic categories. These are: Horizontal Planar Surfaces (HP), which include ground, level terrain, and flat artificial or natural shore platforms; Vertical Planar Surfaces (VP), encompassing the facades of structures like bridges, buildings, and vessels, as well as vertical embankments; Complex Horizontal Terrain (CH), referring to irregular embankments, grasslands and low-lying scrub; and Complex Vertical Structures (CV), representing non-planar, upright natural elements like trees and dense bankside vegetation. Additionally, two further categories are defined for features on and over the water: Open Water Surface (OW) and Under Bridge Clearance Zone (UB), which identify the safe, traversable area beneath bridge structures.
@@
The next step is to perform slope estimation on these voxel sets, as illustrated in Fig. 8. In contrast to methods that derive slope using only boundary heights, our approach calculates the slope through all voxel points within a defined local neighbourhood. This method is more suitable for estimating those structures such as VP, CV, or overhangs.
@@
The parameters $\theta_{\mathrm{ls}}=[a,b,c]^{\top}$ defining the best-fit plane are determined by solving the linear least squares problem shown in Eq. (16). To ensure numerical stability, the solution is obtained by computing the Moore–Penrose pseudoinverse $A^{+}$ of $A$ via singular value decomposition (SVD), rather than by explicitly forming normal equations.
@@
3.4.3 Generation of IENC Compatible Map
@@
0: Shoreline points $\mathcal{S}$, bridge polygons $\mathcal{B}$
@@
8:    $\mathbf{p}\leftarrow\mathbf{p}_{0}+k\cdot s\cdot(\cos\theta,\sin\theta)^{\top}$
@@
32:return  $\mathcal{S},\mathcal{B}$
4 Experiments and Discussion
4.1 Data Collection
@@
Data were collected using a catamaran named the Maverick, with an overall length of 6.10 m and a beam of 2.02 m, as illustrated in Fig. 1. Its draft varies from approximately 0.30 m (lightship) to 0.60 m (fully ballasted). The twin‐hull configuration provides inherent stability, beneficial for sensor data acquisition. The vessel was equipped with the sensor suite detailed in Section 3.2.
We recorded three datasets along this route, as shown in Fig. 9, to cover different typical inland waterway environments and to demonstrate the semantic mapping capability. These trajectories are named as follows, listed in Table 1:
  • 1.
@@
Table 1: Our Inland Waterway Datasets
@@
4.2 Evaluation of Trajectory and Point Cloud Mapping

### v2

@@ III-C1 Preprocessing and Plane Fitting
Given the LiDAR’s non-uniform vertical distribution of its 32 laser beams and a fixed horizontal resolution of $0.2^{\circ}$, each scanning cycle produces a 2D range image, $I_{k}$, of dimension $32\times 1800$ (rows $\times$ columns). Each row corresponds to an individual laser beam, $v_{j}$ represents the scan line index (row), while each column represents sequential horizontal angle increments, $u_{j}$ represents the horizontal angle index (column).
@@
Our system is designed for short-range localization and mapping; the curvature of the Earth can be neglected, and the water surface can be approximated as a plane [48]. Moreover, considering the limitations discussed above posed by the incidence angles of grazing on the surface of the water, the candidate points for the estimation of the plane are restricted to a predefined subset of lower LiDAR beams ($N_{\text{lines}}$), determined by the installation height of the LiDAR installation and the draft of the vessel.
@@
$\mathbf{n}_{s}^{\mathsf{T}}\mathbf{p}+d_{s}=0$ (2)
where $\mathbf{n}_{s}=[a_{s},b_{s},c_{s}]^{\mathsf{T}}\in S^{2}$ is the normal vector of the plane and $d_{s}\in\mathbb{R}$ is its scalar offset from the sensor origin. Subsequently, any point $\mathbf{p}_{j}\in\mathbf{P}_{k}^{valid}$ is classified as a water surface point, setting the corresponding $M_{k}^{surface}(v_{j},u_{j})=1$, if its perpendicular distance to this estimated plane, calculated as $|\mathbf{n}_{s}^{\mathsf{T}}\mathbf{p}_{j}+d_{s}|$, is within a predefined tolerance $\tau_{s}$.
@@
All points belonging to a connected component found via BFS are assigned a unique positive integer identifier in Channel 3 ($M_{k}^{\text{label}}$). To ensure the robustness of subsequent feature extraction, a final filtering step is applied: segments comprising fewer than 30 points are considered insignificant or noise. Points belonging to these small segments are relabeled in Channel 3 as outliers.
III-C2 Feature Extraction
@@
Traditional approaches often extract features based on local range differences. For instance, a common metric (Equation 3, where $|L|$ is the size of the local neighborhood, e.g., 10 points, and $r_{i}$ is the range of point $i$) is:
$c=\frac{1}{|L|\cdot\|r_{i}\|}\left\|\sum_{j\in L,j\neq i}(r_{j}-r_{i})\right\|$ (3)
While this method is computationally efficient and applicable to all directions, it demonstrates limitations in inland waterway environments. In areas with dense vegetation, grass, or gravel, the local surface exhibits numerous small, ”edge-like” bumps. When applying Equation 3, the positive and negative range residuals $(r_{j}-r_{i})$ from neighboring points $p_{j}$ on either side of the central point $p_{i}$ tend to cancel each other out. Consequently, such curvature estimations often overlook subtle surface textures and primarily respond only to more extreme geometric features like significant peaks or occlusions.
To overcome this limitation and effectively capture local surface roughness, thereby distinguishing between smooth and textured surfaces, we employ a variance-based approach to quantify local range variations as detailed in Algorithm  1. Figure 4 illustrates the comparative performance of feature extraction between LOAM-based methods and our optimized approach.
@@
III-C3 LiDAR Odometry, Mapping and Joint Optimization
Our LiDAR odometry and mapping parts build upon LeGO-LOAM [16], with specific adaptations in Joint Optimization. The system employs a two-stage registration process for pose estimation, integrating water surface constraints. The overall factor graph structure is illustrated in Figure 5.
Two-Stage LiDAR Odometry: a two-step Levenberg-Marquardt optimization to estimate the transformation between consecutive scans. First, the planar features $\mathcal{F}_{p}^{t}$ extracted from the water surface and other planar environmental elements are matched to their correspondences using a point-to-plane association in $\mathcal{F}_{p}^{t-1}$ to determine $t_{z},\theta_{roll},\theta_{pitch}]$. Then, the remaining transformation parameters $t_{x},t_{y},\theta_{yaw}]$ are estimated using edge features $\mathcal{F}_{e}^{t}$ matched against $\mathcal{F}_{e}^{t-1}$ through a point-to-line association [15], while applying the previously computed vertical and angular parameters as constraints, resulting in the full 6 DoF transformation. This decoupled registration method significantly reduces computation time and is well-suited to our edge computing scenario.
@@ III-C3 LiDAR Odometry, Mapping and Joint Optimization
Joint Optimization with Water Surface Constraints: The water surface provides a globally reference plane, ideal for mitigating drift that accumulates over long trajectories. Unlike the fragmented planar features typically found in terrestrial environments, the vast and continuous nature of the water surface offers a simple but effective constraint. To take advantage of this, we introduce the water surface constraint into the factor graph optimization framework. We parameterize the water surface as a globally consistent plane $P_{w}$ with parameters $\{\mathbf{n}_{w},d_{w}\}$ as defined in Equation 2.
@@
In this formulation, $\|r\|^{2}_{\Sigma}$ denotes the squared Mahalanobis distance $r^{\mathsf{T}}\Sigma^{-1}r$, where $\Sigma$ is the covariance matrix of the corresponding residual $r$. The term $r_{0}$ represents a combined prior residual on initial knowledge about the system, such as the initial pose $T_{0}$ or an initial estimate for the plane, with $\Sigma_{0}$ as its covariance. The odometry residual, $r_{\mathrm{odom},k}$, for the relative motion $\Delta\widetilde{T}_{k,k+1}$ measured between consecutive poses $T_{k}$ and $T_{k+1}$, is given by
@@
where $\mathrm{Log}(\cdot)$ maps an $SE(3)$ element to its tangent space representation. Similarly, the loop closure residual $r_{\mathrm{loop},ij}$, constrains the measured relative transform $\Delta\widetilde{T}_{i,j}$ between non-consecutive poses $T_{i}$ and $T_{j}$. This transform is computed using ICP [52] to verify a potential match between keyframes, and the resulting constraint is added to a pose graph optimized by iSAM2 [42]. Note that only the experiment in area A uses this technique. The residual (where $(i,j)\in\mathcal{L}$ denotes the set of loop closures) is defined as
@@
$(R_{m}n_{s,m})^{\mathsf{T}}x_{w}+\bigl(d_{s,m}-(R_{m}n_{s,m})^{\mathsf{T}}t_{m}\bigr)=0,$ (9)
@@
$\begin{split}r_{\mathrm{plane},m}(T_{m},P_{w})&=P_{w}\ominus\widehat{P}_{w,m}\\ &=\begin{pmatrix}e_{N}\!\bigl(n_{w},\hat{n}_{w,m}\bigr)\\ d_{w}-\hat{d}_{w,m}\end{pmatrix}\\ &=\begin{pmatrix}e_{N}\!\bigl(n_{w},R_{m}n_{s,m}\bigr)\\ d_{w}-\bigl(d_{s,m}-(R_{m}n_{s,m})^{\mskip 2.0mu\mathsf{T}}t_{m}\bigr)\end{pmatrix}\in\mathbb{R}^{3}.\end{split}$ (11)
@@
$\hat{\xi}=-\frac{\arccos(n_{w}^{\mathsf{T}}\hat{n}_{w,m})}{\sqrt{1-(n_{w}^{\mathsf{T}}\hat{n}_{w,m})^{2}}}\,(\hat{n}_{w,m}-(n_{w}^{\mathsf{T}}\hat{n}_{w,m})\,n_{w}),$ (12)
@@
By leveraging the water surface as a global geometric constraint, our approach reduces vertical drift, as visualized in Figure 6.
@@
III-D Structural Semantic Map Conversion
III-D1 Semantic voxel transformation
In the preceding section, a map comprising structured point cloud features was generated. Each point in this cloud has been classified with semantic labels, including the water surface, flat points, or rough points. To efficiently store and query this semantic information within the global map, the ”dense” semantically annotated point cloud is transformed into a ”sparse” voxel grid representation. For each voxel, a probabilistic estimate of its semantic terrain class is maintained. This is achieved by adapting the log-odds update mechanism, commonly used in probabilistic occupancy grids [44], to instead fuse evidence over time for competing semantic labels.
The workspace is discretized into a 3D voxel grid with a uniform resolution $r$, set to $0.15\text{ m}$ in this work. This resolution is chosen considering a balance between memory consumption, computational efficiency, and the typical scale of objects of interest in inland waterways. The entire transformation process is depicted in Figure 7. For each voxel $v_{i}$, we maintain a hash map to store its semantic state $S(v_{i})$. This state involves accumulating evidence for the semantic classes: water surface, flat terrain, and rough terrain. The ’unknown’ category is assigned if a voxel has insufficient evidence for any specific class.
@@ III-D1 Semantic voxel transformation
Points pre-classified as water surface by the LiDAR odometry module are handled directly, if such a point falls into voxel $v_{i}$, the voxel is immediately assigned the ”water” semantic label with high confidence.
@@
$S(v_{i})=\begin{cases}\text{"flat"}&\text{if }P_{i}^{\text{flat}}-0.5>\tau_{\text{flat}}\\ \text{"rough"}&\text{if }P_{i}^{\text{flat}}-0.5<-\tau_{\text{rough}}\\ \text{"unknown"}&\text{otherwise}\end{cases}$
III-D2 2D structural semantic map construction
The 3D semantic voxel map is subsequently projected onto a unified 2D structural semantic map. While it refers on the conversion method [41], this variant targets USV rather than UGV/UAV scenarios and adopts different map architectures for refined slope estimation and richer semantic inference. To effectively map the complex environment encountered by USVs in inland waterways, this framework broadly classifies non water surface areas into four structural semantic categories. These are: Horizontal Planar Surfaces (HP), which include water surface, level terrain, and flat artificial or natural shore platforms; Vertical Planar Surfaces (VP), encompassing the facades of structures like bridges, buildings, and vessels, as well as vertical embankments; Complex Horizontal Terrain (CH), referring to irregular embankments, grasslands and low-lying scrub; and Complex Vertical Structures (CV), representing non-planar, upright natural elements like trees and dense bankside vegetation. Additionally, two further categories are defined for features on and over the water: Open Water Surface (OW) and Under Bridge Clearance Zone (UB), which identify the safe, traversable area beneath bridge structures.
@@
The next step is to perform slope estimation on these voxel sets, as illustrated in Figure 7. In contrast to methods that derive slope using only boundary heights, our approach calculates the slope through all voxel points within a defined local neighbourhood. This method is more suitable for estimating those structures such as VP, CV, or overhangs.
@@
The parameters $\theta_{\mathrm{ls}}=[a,b,c]^{\mathsf{T}}$ defining the best-fit plane are determined by solving the linear least squares problem shown in Eq. (16). To ensure numerical stability, the solution is obtained by computing the Moore–Penrose pseudoinverse $A^{+}$ of $A$ via singular value decomposition (SVD), rather than by explicitly forming normal equations.
@@
III-D3 Generation of IENC Compatible Map
@@
0: Shoreline points $\mathcal{S}_{\text{shore}}$, bridge polygons $\mathcal{B}$
@@
8:    $\mathbf{p}\leftarrow\mathbf{p}_{0}+k\cdot s\cdot(\cos\theta,\sin\theta)^{\mathsf{T}}$
@@
32:return $\mathcal{S}_{\text{shore}},\mathcal{B}$
IV Experiments and Discussion
IV-A Data Collection
@@
Data were collected using a catamaran named the Maverick, with an overall length of 6.10 m and a beam of 2.02 m, as illustrated in Figure 1. Its draft varies from approximately 0.30 m (lightship) to 0.60 m (fully ballasted). The twin‐hull configuration provides inherent stability, beneficial for sensor data acquisition. The vessel was equipped with the sensor suite detailed in Section III-B.
We recorded three datasets along this route, as shown in Figure 8, to cover different typical inland waterway environments and to demonstrate the semantic mapping capability. These trajectories are named as follows, listed in Table I:
@@
TABLE I: Our Inland Waterway Datasets
@@
IV-B Evaluation of Trajectory and Point Cloud Mapping